The total charge entering a terminal is q=(10-10e-2t) mC. Calculate current at t=0.5 s


1
Expert's answer
2022-02-14T21:39:03-0500

i=dq/dt

= d((10-10e-2t) )/dt

= 10*2e-2t

I(t=0.5 sec)= 20e-1